OpenStudy (anonymous):

Two altitudes of a triangle have lengths 12 and 14. What is the longest possible length of the third altitude, if it is a positive integer? I have no idea what to do. Someone please help!

OpenStudy (anonymous):

Let the sides of the triangle be a, b, and c, and let the altitudes to these sides be h_A = 12, h_B = 14, and h_C, respectively. Let K denote the area of the triangle. Then 1/2 ah_A = 1/2 bh_B = 1/2 ch_C = K so a = 2K/h_A, b = 2K/h_B, c = 2K/h_C By the triangle inequality, a + b > c a + c > b b + c > a Substituting our formulas above, we get 2K/h_A + 2K/h_B > 2K/h_C 2K/h_A + 2K/h_C > 2K/h_B 2K/h_B + 2K/h_C > 2K/h_A which simplify to 1/h_A + 1/h_B > 1/h_C 1/h_A + 1/h_C > 1/h_B 1/h_B + 1/h_C > 1/h_A Substituting h_A = 12 and h_B = 14, we get 1/12 + 1/14 > 1/h_C 1/12 + 1/h_C > /14 1/14 + 1/h_C > 1/12 From the first inequality, 1/h_C < 1/12+1/14=13/84 so h_C > 84/13. Since 1/12 > 1/14, the second inequality is always satisfied. And from the third inequality, 1/h_C >1/12 - 1/14 = 1/84 so h_C < 84. Therefore, the largest possible value of h_C is 83.

OpenStudy (anonymous):

Thanks!